OpenStudy (anonymous):

Tom and John are to sit in a row of 8 seats with 6 other people. How many ways are there for the 8 people to sit if Tom sits to the left of John (but not necessarily beside him) ?

OpenStudy (astrophysics):

<3

OpenStudy (astrophysics):

It's a protest to get your 100 back sorry @Draven_03

OpenStudy (astrophysics):

Lets do this problem though

OpenStudy (astrophysics):

Ok so

OpenStudy (anonymous):

lol, thanks though

OpenStudy (astrophysics):

Err yeah permutations

OpenStudy (anonymous):

number the seats one to 8 from left to right john can't sit in seat 1, but he can sit in seat two

OpenStudy (anonymous):

if he does there is one spot for tom, and oh crap there are 7 people not 8 !!

OpenStudy (anonymous):

oh no 8 good

OpenStudy (anonymous):

if john sits in seat 2 there are 6! ways to fill the other chair this is really going to be a drag

OpenStudy (anonymous):

if john sits in seat 3 there are 2 choices for tom and 6 choices for the other seat and 5! choices for the remaining i am probably not doing this the best way, i hate counting, but it john sits in seat 3 i get \[2\times 6!\times 5!\]

OpenStudy (anonymous):

no that is wrong \[2\times 6\times 5!\] or better yet \(2\times 6!\)

OpenStudy (anonymous):

i think does that make sense if john is in seat 3?

OpenStudy (anonymous):

tbh im lost now, are all of those answers equal to each other?

OpenStudy (anonymous):

i don't know lets go slow

OpenStudy (anonymous):

thank you for all of your help

OpenStudy (anonymous):

i don't think so

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

A john sits in seat 2 tom must sit in seat 1, then 6! for the remaining chairs so far so good?

OpenStudy (anonymous):



OpenStudy (anonymous):

good so far? just fyi i hate this stuff, so if someone has a snazzier way i would be happy to shut up and see it

OpenStudy (astrophysics):

It's your day today sat go ahead!

OpenStudy (anonymous):

lol @Draven_03 is it clear or not clear so far?

OpenStudy (anonymous):

actually i don't think it is going to turn out that bad

OpenStudy (anonymous):

lol sorry, my screen froze

OpenStudy (sweetburger):

thats a lot of bikes

OpenStudy (anonymous):

but yea that makes sense

OpenStudy (anonymous):

yeah i think mine did too i got an answer but lets go slow (especially since it might not be right) so back to the beginning, john sits in seat 2, then 6! ways to fill up the other chairs

OpenStudy (astrophysics):

Yes, we want lance armstrong to be recognized for winning his medals on wikipedia

Miracrown (miracrown):

LOL

OpenStudy (anonymous):

john sits in seat 3, then two choices for tom, 6! for the other chairs, so \(2\times 6!\) for seat 3

OpenStudy (anonymous):

damn where'd i put my dope ... ok seat 4 quick 2 choices for tom, 6! for other seats \(3\times 6!\) pattern is emerging

OpenStudy (anonymous):

i meant 3 choices for tom, etc

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

i am not willing to bet more that $14 that this answer is correct, but it looks like to me \[(1+2+3+4+5+6+7)\times 6!\]

OpenStudy (anonymous):

oh lord is that really just \[\frac{8!}{2}\]??"

OpenStudy (astrophysics):

Didn't i say that earlier

OpenStudy (astrophysics):

!!

OpenStudy (anonymous):

all right all you bikes ....

OpenStudy (anonymous):

why didn't you say so earlier @Astrophysics ?

OpenStudy (astrophysics):

The answer seemed insane to me though idk, I did but I erased it

OpenStudy (astrophysics):

deleted

OpenStudy (anonymous):

lol, I could have mentioned to you that it was actually a timed question, and we ran out of time a long time ago. ps i didn't even know it was timed until i was out of time lol

OpenStudy (anonymous):

how did you know it so quick and why did you let me struggle so long?

OpenStudy (anonymous):

half left and half right doe

OpenStudy (astrophysics):

Yup lol, and it's fun

OpenStudy (anonymous):

i am going to go watch perry mason told you i hate these, especially when there is some complete trivial explanation you (I) don't see

OpenStudy (anonymous):

I thought your examples would become clear, and i would learn how to do it; which in a way i suppose i kind of did

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

sheer reason is always better than computation

OpenStudy (anonymous):

whats perry mason?

OpenStudy (anonymous):

what's perry mason!!!???

OpenStudy (astrophysics):

I wonder if Perry Mason could've handled this question as easily mhmm

OpenStudy (anonymous):

i am sure della could have

OpenStudy (anonymous):

lol, isn't perry mason a tv show ? or a person on a tv show ?

OpenStudy (anonymous):

it is a very very long running tv show from the fifties to the early sixties based on novels by erle stanley garner

candycove (candycove):

LOOOLLLL

OpenStudy (astrophysics):



Miracrown (miracrown):

Happy satellite day errrbody! <3

Zeronknight (zeronknight):

Um, what's with the satellites for???

Hayhayz (hayhayz):

Happy satellite day ☺

OpenStudy (astrophysics):

We can make a problem out of this 3 mod cyclists are to bike in a row...yeah ok

Miracrown (miracrown):

Today, the first ever satellite was invented! :^)

Zeronknight (zeronknight):

^Learn something new everyday c:

OpenStudy (anonymous):

there are a lot of bikes

OpenStudy (♪chibiterasu):

January 29, 2016. Satellite memorial day.

OpenStudy (♪chibiterasu):

An unfair loss.

Zeronknight (zeronknight):

LOL, why did he lose his 100 though?

candycove (candycove):

rip

OpenStudy (♪chibiterasu):

Because a certain sometime did a certain something and now he has to pay for it rofl.